OpenStudy (wesdg1978):

f(x)=5/x^2+sinx How do you find f '(x)?

OpenStudy (anonymous):

is it \[f(x)=\frac{5}{x^2}+\sin(x)\] or \[f(x)=\frac{5}{x^2+\sin(x)}\]?

OpenStudy (wesdg1978):

The first one

OpenStudy (anonymous):

ok then take the derivative piece by piece you know the derivative of sine?

OpenStudy (wesdg1978):

cos

OpenStudy (anonymous):

ok good how about the derivative of \(\frac{5}{x^2}\)?

OpenStudy (wesdg1978):

-10x/x4

OpenStudy (anonymous):

if you want you can write is as \[5x^{-2}\] and use the power rule

OpenStudy (wesdg1978):

-10x^-3

OpenStudy (anonymous):

or you can do what you did, but then make sure to cancel an \(x\) top and bottom, because \[-\frac{10x}{x^4}=-\frac{10}{x^3}\]

OpenStudy (anonymous):

same answer either way, and now you are done answer : \[-\frac{10}{x^3}+\cos(x)\]

OpenStudy (wesdg1978):

Thanks

OpenStudy (anonymous):

yw